OpenStudy (anonymous):

The roof of a schoolhouse is shaped like an isosceles triangle. The base of the triangle measures 8 meters and one side measures 5 meters. Find the altitude.

OpenStudy (anonymous):

3

OpenStudy (saifoo.khan):

3m

OpenStudy (anonymous):

The roof is a composite of two 3,4,5 triangles. In Pythagorean theorem, 3, 4, 5 triangles work out to whole numbers due to the Pythagorean theorem which is a^2 = b^2 + c^2 or 5^2 = 4^2 + 3^2 or 25 = 16 + 9. the altitude, or height of the triangles and therefore the roof turns out to be 3 m. Note: Diagram shows cm, but just think of them as m, lol